f={(1,2),(2,3),(3,3)},g={(1,0)(2,3),(3,2),(4,8)}
Find f/g=

Answer by greenestamps(13196)



f/g can only be defined for x value for which both f and g are defined: {1,2,3}.

However, (f/g)(1) = 2/0 is undefined; so 1 is not in the domain of f/g.

(f/g)(2) = f(2)/g(2) = 3/3 = 1
(f/g)(3) = f(3)/g(3) = 3/2

ANSWER: f/g = {(2,1),(3,3/2)}